A 67.0 kg person jumps from rest off a 3.00 m high towerstraight down into the water. Neglect air resistance duringthe descent. She comes to rest 1.10 m under the surface ofthe water. Determine the magnitude of the average force that thewater exerts on the diver. This force isnonconservative.
Answer #1
You can do this as work-energy:
   total work   = ΔK    but since she started and ended at rest,initial and final K are both zero!
So    total work =0      what forces did work?
Gravity:     force * distance = mgΔh =   67 * 9.8 * 4.10 = 2692 Joules
Notice the total distance she travels... 3.00 +1.10   because gravity still acts on her even when she isin the water!
Now the work by the water:     - Force * dist =    - F *1.10    where it is negative because the forcepushes up on her, slowing her down, while she travels down. Whenforce and distance are in opposite directions, work isnegative.
Now:     Work by gravity + work bywater = 0
               2692      -   F *1.10   = 0       or      2692 = F *1.10    or   F = 2447 Newtons!
Answer #2
let initial velocity at the surface of water be u
as the person falls from height h = 3m
u = √2gh = √2*9.8*3 = √58.8   =7.668 m/s2
final velocity of the person inside water v =0
distance travelled inside water s = 1.1m
apply v2- u2 = -2as
plug the values and solve for a (retardation)
mass of person m = 67kg
force = ma = 67*a N
plug the value of a